# COALA: Numerically Stable and Efficient Framework for Context-Aware Low-Rank Approximation

### v1

@@ Appendix E Convergence Proofs (Without the Full-Rank Condition)
Before proving the theorem, we first establish the following auxiliary lemma.
Lemma 3.
Let $W\in\mathbb{R}^{m\times n},B\in\mathbb{R}^{n\times n}$ be an arbitrary matrices. Define
$A=W\begin{bmatrix}\Lambda&0\\ 0&0\end{bmatrix},$
$\hat{A}=W\begin{bmatrix}\Lambda+B&0\\ 0&\sqrt{\varepsilon}I\end{bmatrix},$
where $\Lambda=\operatorname{diag}(\lambda_{1},\dots,\lambda_{k})$ with $k\geq r$. Let $U_{r}$ and $\hat{U}_{r}$ be the first $r$ left singular vectors of $A$ and $\hat{A}$, respectively, and suppose $\sigma_{r}(A)\neq\sigma_{r+1}(A)$. Then it holds that
$\|U_{r}U_{r}^{\top}-\hat{U}_{r}\hat{U}_{r}^{\top}\|_{2}\leq\frac{2\|W\|_{2}^{2}\bigl{(}2\lambda_{\max}\|B\|_{2}+\max(\|B\|^{2}_{2},\varepsilon)\bigr{)}}{\sigma_{r}(A)^{2}-\sigma_{r+1}(A)^{2}}.$
Proof.
Let
$E=A-\hat{A}=W\begin{bmatrix}B&0\\ 0&\sqrt{\varepsilon}I\end{bmatrix}.$
Consider
$\hat{A}\hat{A}^{\top}=(A+E)(A+E)^{\top}=AA^{\top}+AE^{\top}+EA^{\top}+EE^{\top}.$
Then
$AA^{\top}-\hat{A}\hat{A}^{\top}$ $=AE^{\top}+EA^{\top}+EE^{\top}=$
$=W\begin{bmatrix}\Lambda&0\\ 0&0\end{bmatrix}\begin{bmatrix}B^{\top}&0\\ 0&\varepsilon I\end{bmatrix}W^{\top}+W\begin{bmatrix}B&0\\ 0&\varepsilon I\end{bmatrix}\begin{bmatrix}\Lambda&0\\ 0&0\end{bmatrix}W^{\top}+W\begin{bmatrix}BB^{\top}&0\\ 0&\varepsilon I\end{bmatrix}W^{\top}=$
$=W\left(\begin{bmatrix}\Lambda B^{\top}&0\\ 0&0\end{bmatrix}+\begin{bmatrix}B\Lambda&0\\ 0&0\end{bmatrix}+\begin{bmatrix}BB^{\top}&0\\ 0&\varepsilon I\end{bmatrix}\right)W^{\top}.$
Hence,
$\|AA^{\top}-\hat{A}\hat{A}^{\top}\|_{2}$ $=\|AE^{\top}+EA^{\top}+EE^{\top}\|_{2}=$
$=\left\|W\left(\begin{bmatrix}\Lambda B^{\top}&0\\ 0&0\end{bmatrix}+\begin{bmatrix}B\Lambda&0\\ 0&0\end{bmatrix}+\begin{bmatrix}BB^{\top}&0\\ 0&\varepsilon I\end{bmatrix}\right)W^{\top}\right\|_{2}\leq$
$\leq\|W\|^{2}_{2}\left\|\begin{bmatrix}\Lambda B^{\top}&0\\ 0&0\end{bmatrix}+\begin{bmatrix}B\Lambda&0\\ 0&0\end{bmatrix}+\begin{bmatrix}BB^{\top}&0\\ 0&\varepsilon I\end{bmatrix}\right\|_{2}\leq$
$\leq\|W\|_{2}^{2}\left(\left\|\begin{bmatrix}\Lambda B^{\top}&0\\ 0&0\end{bmatrix}\right\|_{2}+\left\|\begin{bmatrix}B\Lambda&0\\ 0&0\end{bmatrix}\right\|_{2}+\left\|\begin{bmatrix}BB^{\top}&0\\ 0&\varepsilon I\end{bmatrix}\right\|_{2}\right)\leq$
$\leq\|W\|_{2}^{2}\left(\|\Lambda B^{\top}\|_{2}+\|B\Lambda\|_{2}+\max(\|BB^{\top}\|_{2},\varepsilon)\right)\leq$
$\leq\|W\|_{2}^{2}\left(\|\Lambda\|_{2}\|B^{\top}\|_{2}+\|B\|_{2}\|\Lambda\|_{2}+\max(\|B\|_{2}^{2},\varepsilon)\right)\leq$
$\leq\|W\|_{2}^{2}\bigl{(}2\lambda_{\max}\|B\|_{2}+\max(\|B\|_{2}^{2},\varepsilon)\bigr{)}.$
Denote by $U_{r}$ and $\hat{U}_{r}$ the left singular vectors of $AA^{\top}$ and $\hat{A}\hat{A}^{\top}$, respectively. Applying the Davis–Kahan Theorem 4 we obtain
$\|U_{r}U_{r}^{\top}-\hat{U}_{r}\hat{U}_{r}^{\top}\|_{2}$ $\leq\frac{2\|AA^{\top}-\hat{A}\hat{A}^{\top}\|_{2}}{\sigma_{r}(AA^{\top})-\sigma_{r+1}(AA^{\top})}=$
$=\frac{2\|AA^{\top}-\hat{A}\hat{A}^{\top}\|_{2}}{\sigma_{r}(A)^{2}-\sigma_{r+1}(A)^{2}}\leq$
$\leq\frac{2\|W\|_{2}^{2}\bigl{(}2\lambda_{\max}\|B\|_{2}+\max(\|B\|_{2}^{2},\,\varepsilon)\bigr{)}}{\sigma_{r}(A)^{2}-\sigma_{r+1}(A)^{2}}$
Finally, we observe that the left singular vectors of $A$ and $AA^{\top}$ coincide, and likewise for $\hat{A}$ and $\hat{A}\hat{A}^{\top}$. Consequently, the derived estimate is precisely the solution of interest. ∎
@@
Let us consider the unitary eigenvalue decomposition $P\Lambda P^{\top}=XX^{\top}$, where $P$ is unitary and $\Lambda$ consists of non-negative values.
From Propositions 1 and 3 , we have:
$W_{0}=U_{r}U_{r}^{\top}W,$
$W_{\mu}=V_{r}V_{r}^{\top}W,$
where $U_{r}$ and $V_{r}$ are the top $r$ left singular vectors of $WP\Lambda^{1/2}P^{\top}$ and $WP(\Lambda+\mu I)^{1/2}P^{\top}$, respectively.
Let $\widetilde{W}=WP$. Note that $U_{r}$ and $V_{r}$ are the top $r$ left singular vectors of the matrices $\widetilde{W}\Lambda^{1/2}$ and $\widetilde{W}(\Lambda+\mu I)^{1/2}$, respectively, since right multiplication by an orthogonal matrix does not affect the left singular vectors.
Observe that $X^{\top}=Q(P\Lambda^{1/2}P^{\top})$ is the polar decomposition of $X$ for some matrix $Q$ with orthogonal columns. Then we have
$WX=WP\Lambda^{1/2}P^{\top}Q^{\top}=\widetilde{W}\Lambda^{1/2}(P^{\top}Q^{\top}).$
Since $P^{\top}Q^{\top}$ consists of orthogonal rows, the sets of singular values of the matrices $WX$ and $\widetilde{W}\Lambda^{1/2}$ coincide. Therefore,
$\sigma_{r}(\widetilde{W}\Lambda^{1/2})\neq\sigma_{r+1}(\widetilde{W}\Lambda^{1/2}).$
Assume that $\operatorname{rank}(X)=k\geq r$. Let us define
$\Lambda=\begin{bmatrix}S&0\\ 0&0\end{bmatrix}.$
Then, we set
$A=\widetilde{W}\Lambda^{1/2}=\widetilde{W}\begin{bmatrix}S^{1/2}&0\\ 0&0\end{bmatrix},$
$\hat{A}=\widetilde{W}(\Lambda+\mu I)^{1/2}=\widetilde{W}\begin{bmatrix}S^{1/2}+B&0\\ 0&\sqrt{\mu}I\end{bmatrix},$
where $B=(S+\mu I)^{1/2}-S^{1/2}$. Applying Lemma 3, we obtain
$\|U_{r}U_{r}^{\top}-V_{r}V_{r}^{\top}\|_{2}\leq\frac{2\|\widetilde{W}\|_{2}^{2}\left(2\sqrt{s_{\max}}\|B\|_{2}+\max\left(\mu,\|B\|_{2}^{2}\right)\right)}{\sigma_{r}(\widetilde{W}\Lambda^{1/2})^{2}-\sigma_{r+1}(\widetilde{W}\Lambda^{1/2})^{2}}.$
By Lemma 2, we have $\|B\|_{2}\leq\dfrac{\mu}{2\sqrt{s_{\min}}}=\dfrac{\mu}{2\sigma_{k}(X)}$. Substituting this into the inequality, we get
$\|U_{r}U_{r}^{\top}-V_{r}V_{r}^{\top}\|_{2}\leq\frac{2\|\widetilde{W}\|_{2}^{2}\left(\dfrac{\sqrt{s_{\max}}}{\sigma_{k}(X)}\mu+\max\left(\mu,\dfrac{\mu^{2}}{4\sigma_{k}(X)^{2}}\right)\right)}{\sigma_{r}(\widetilde{W}\Lambda^{1/2})^{2}-\sigma_{r+1}(\widetilde{W}\Lambda^{1/2})^{2}}.$
Furthermore, since $\|\widetilde{W}\|_{2}=\|WP\|_{2}=\|W\|_{2}$ due to the unitary invariance of the spectral norm, and from our earlier observations $\sigma_{i}(\widetilde{W}\Lambda^{1/2})=\sigma_{i}(WX)$, we have
$\|U_{r}U_{r}^{\top}-V_{r}V_{r}^{\top}\|_{2}\leq\frac{2\|W\|_{2}^{2}\left(\dfrac{\sqrt{s_{\max}}}{\sigma_{k}(X)}\mu+\max\left(\mu,\dfrac{\mu^{2}}{4\sigma_{k}(X)^{2}}\right)\right)}{\sigma_{r}(WX)^{2}-\sigma_{r+1}(WX)^{2}}.$
Noting that $s_{\max}=\sigma_{1}(X)^{2}$ and factoring out $\mu$, we conclude
$\|U_{r}U_{r}^{\top}-V_{r}V_{r}^{\top}\|_{2}\leq\frac{2\|W\|_{2}^{2}\left(\dfrac{\sigma_{1}(X)}{\sigma_{k}(X)}+\max\left(1,\dfrac{\mu}{4\sigma_{k}(X)^{2}}\right)\right)}{\sigma_{r}(WX)^{2}-\sigma_{r+1}(WX)^{2}}\cdot\mu.$
@@ Appendix F Implementation Details
Learning Rate $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$

### v2

@@ Appendix E Convergence Proofs (Without the Full-Rank Condition)
By Proposition 1 , we obtain
$W_{0}=U_{0}U_{0}^{\top}W,$
where $U_{0}$ denotes the first $r$ left singular vectors of $WX$, and $H_{0}=(XX^{\top})^{1/2}$. Analogously, using Proposition 3 , we have
$W_{\mu}=U_{\mu}U_{\mu}^{\top}W,$
where $U_{\mu}$ denotes the first $r$ left singular vectors of $WH_{\mu}$, and $H_{\mu}$ = $(XX^{\top}+\mu I)^{1/2}$. However, we can get the matrices $U_{0}$ and $U_{\mu}$ are defined as the matrices of the first $r$ left singular vectors obtained from the singular value decompositions:
$U_{0}\leftarrow\mathrm{SVD}\bigl(WXX^{\top}W^{\top}\bigr),\qquad U_{\mu}\leftarrow\mathrm{SVD}\bigl(W(XX^{\top}+\mu I)W^{\top}\bigr).$
Here we use the fact that the left singular vectors of a matrix $A$ coincide with the eigenvectors vectors (same that left singular in this case) of $AA^{\top}$.
Consider the perturbation of the matrix $WXX^{\top}W^{\top}$:
$\|WXX^{\top}W^{\top}-W(XX^{\top}+\mu I)W^{\top}\|_{2}=\mu\|WW^{\top}\|_{2}=\mu\|W\|_{2}^{2}.$
By Applying Davis-Kahan Theorem 4, we obtain
$\|U_{0}U_{0}^{\top}-U_{\mu}U_{\mu}^{\top}\|_{2}\leq\frac{2\|WXX^{\top}W^{\top}-W(XX^{\top}+\mu I)W^{\top}\|_{2}}{\sigma_{r}(WXX^{\top}W^{\top})-\sigma_{r+1}(WXX^{\top}W^{\top})}.$
Since $\sigma_{k}(WXX^{\top}W^{\top})=\sigma_{k}^{2}(WX)$, this yields
$\|U_{0}U_{0}^{\top}-U_{\mu}U_{\mu}^{\top}\|_{2}\leq\frac{2\|W\|_{2}^{2}}{\sigma_{r}^{2}(WX)-\sigma_{r+1}^{2}(WX)}\,\mu.$
Combining this bound with $W_{0}=U_{0}U_{0}^{\top}W$ and $W_{\mu}=U_{\mu}U_{\mu}^{\top}W$, we arrive at
$\|W_{0}-W_{\mu}\|_{F}$ $=\|(U_{0}U_{0}^{\top}-U_{\mu}U_{\mu}^{\top})W\|_{F}$
$\leq\|U_{0}U_{0}^{\top}-U_{\mu}U_{\mu}^{\top}\|_{2}\,\|W\|_{F}$
$\leq\frac{2\|W\|_{2}^{2}\|W\|_{F}}{\sigma_{r}^{2}(WX)-\sigma_{r+1}^{2}(WX)}\,\mu.$
@@ Appendix F Implementation Details
Learning Rate $\times 10^{-4} 1×1$$\times 10^{-4}$ $\times 10^{-4} 1×1$$\times 10^{-4}$